TC=50+10Q+0.04Q^(2 )
MC=10+0.08Q
TR=40+20Q+0.02Q^2
MR=20+0.04Q
Based on the above information
Find the quantity that th firm produces
Find the price at which the firm produces
Find the total cost, total variable cost and total fixed cost
How much does the firm need to earn to avoid from closing its operations
1
Expert's answer
2019-10-25T10:50:33-0400

Quantity that the firm produces 

250 units

"MC=MR"


"10+0.08Q=20+0.04Q"

"Q=250"

Price at which the firm produces 

price 30$

"MR=" "20+0.04Q"

"20+0.04*(250)"

"P=30"

Total cost, total variable cost and total fixed cost

Total cost $5050

"TC=50+10Q+0.04Q^2"

"TC=50+10*(250)+0.04(250)^2"

"TC=5050"

Fixed cost $50

"TC=50+(10*0)+0.04(0)^2"

"FC=50"

Variable cost $5000

"VC=TC-FC"

"VC=5050-50"

"VC=5000"

Avoid from closing its operations

"P=AVC"


"AVC=5000\/250"

"AVC=20" If the price below down $20 firm must be shutdown


"AVC=MC"

"20=10+0.08Q"

"Q=125"


"TR=40+20Q+0.02Q^2"

"TR=40+20*125+0.02*125^2"

"TR=2852.5"

 Earn $2852.5 to avoid from closing operations